# LayerPeeler: Autoregressive Peeling for Layer-wise Image Vectorization

### v1

@@ Appendix B More Results
For better accessibility and visualization, we have created a static project website here. The website contains:
@@
Appendix C System Prompts
This section lists the system prompts used by the LayerPeeler system for various Vision-Language Model tasks, covering initial layer graph construction (SectionC.1), layer graph maintenance (Section C.2), instance prompt and bounding box generation (Section C.3), isolated layer annotation (Section C.4), and the specific prompt for the ablation study (Section C.5).
C.1. Initial Layer Graph Construction
@@
C.2. Layer Graph Update
@@
C.3. Instance Mask, Box, and Prompt Generation
@@
C.4. Isolated Layer Annotation
@@
C.5. Ablation Study: Direct Top Layer Detection

### v2

@@ Appendix G More Results
For better accessibility and visualization, we have created a static project website in the supplementary material. The website contains:
@@
Appendix H System Prompts
This section lists the system prompts used by the LayerPeeler system for various Vision-Language Model tasks, covering initial layer graph construction (SectionH.1), layer graph maintenance (Section H.2), instance prompt and bounding box generation (Section H.3), isolated layer annotation (Section H.4), and the specific prompt for the ablation study (Section H.5).
H.1. Initial Layer Graph Construction
@@
H.2. Layer Graph Update
@@
H.3. Instance Mask, Box, and Prompt Generation
@@
H.4. Isolated Layer Annotation
@@
H.5. Ablation Study: Direct Top Layer Detection